OpenStudy (anonymous):

The area of the region bounded above by y=1+sec^2x , below by y=0 on the left by x=0 and on the right by x=pi/4 is approximately... a) 1 b) 1.25 c) 1.5 d) 1.75 e)2

OpenStudy (accessdenied):

Have you tried anything so far on this problem that you could not get to work? Or are you not sure how to begin?

OpenStudy (anonymous):

i am not sure how to begin

OpenStudy (accessdenied):

Well, what we should be looking for is an integration expression over our region that sums up all the areas from x=0 to x=pi/4 under the y=1+sec^2 x graph and above the y=0 graph. A rough sketch of the graph: 

OpenStudy (accessdenied):

 We should add up all these rectangular areas y * dx = (1 + sec^2 ) dx, over the region x=0 to x=pi/4. That is the definition of our definite integral: \( \displaystyle \int_{0}^{\pi/4} (1 + \sec^2 x) \ dx \) Is this setup clear?